OpenStudy (anonymous):

PLEASE HELP! MEDAL WILL BE REWARDED! Write the sum using summation notation, assuming the suggested pattern continues. 100 + 121 + 144 + 169 + ... + n2 + ... Just need to make sure i have the correct answer...

OpenStudy (anonymous):

\[\sum_{n=10}^{n}n^2\] this is my answer! Is this correct?

OpenStudy (sirm3d):

since there is no last term in your sum, there is no need to put any upper limit.

OpenStudy (anonymous):

So everything else is correct just no upper limit? @sirm3d

OpenStudy (sirm3d):

yes, @CarlyLiberty

OpenStudy (anonymous):

100+21=121+21=142+2=144+21=165+3=169